OpenStudy (anonymous):

Starting with the data values 70 and 100, add three data values to the sample so that the mean is 86, the median is 82, and the mode is 82. Please show all of your work.

OpenStudy (lgbasallote):

have you tried trial and error?

OpenStudy (anonymous):

yes it just confuses me a bit more

OpenStudy (lgbasallote):

actually....it's not much of "error"... you know the start is 70 and the last is 100...then there are three blanks in the middle 70 _ _ _ 100 right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then it says the median is 82...so what do you think should you do?

OpenStudy (anonymous):

theoretically, it should look something like 70+82+82+x+100/5= 66.8

OpenStudy (lgbasallote):

exactly

OpenStudy (lgbasallote):

except... it says the mean is 86 so it should be (72 + 82 + 82 + x + 100)/5 = 86

OpenStudy (anonymous):

ohh. So, then how do you find that x?

OpenStudy (lgbasallote):

that should be \[\frac{70 + 82 + 82 + x + 100}5 = 86\]

OpenStudy (lgbasallote):

to solve for x, you should multiply both sides by 5 first

OpenStudy (anonymous):

so have 5*86, which comes to 430, and I believe we can subtract the original 334 by 430, right?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

awesome. That missing number would come to be 96. Thank you!

OpenStudy (lgbasallote):

right